# Stand-In: A Lightweight and Plug-and-Play Identity Control for Video Generation

### v3

@@ 1 Introduction
With the rapid advancement of diffusion models [NEURIPS2020_4c5bcfec, podell2023sdxl, Peebles_2023_ICCV], video generation [opensora, opensora2, kong2024hunyuanvideo, hong2022cogvideo] has become a pivotal aspect of generative AI. Among its diverse applications, identity-preserving video generation holds profound significance. The goal of this task is to generate high-quality videos that consistently maintain the identity of a given reference image containing a face. It has widespread utility across film, advertising and gaming industries, etc.
Existing methods can be roughly classified into two categories. Early methods [ID-Animator, consistid] use an explicit face encoder for identity feature extraction, while recent methods [hu2025hunyuancustom, liu2025phantom] fully train the diffusion transformer. However, face-encoder-based methods lack flexibility and struggle to capture fine facial details essential for high-quality video generation. The full-parameter training methods require huge training resources and lack compatibility with other applications. Achieving robust identity preservation in a lightweight way remains critical yet challenging.
To overcome these limitations, we leverage the pre-trained VAE from the video generation model itself, enabling the conditional image to be mapped directly into the same latent space as the video. This approach naturally utilizes the model’s inherent capabilities to extract rich and detailed facial features, offering a more integrated and effective solution. Specifically, we employ restricted self-attention with conditional position mapping to merge the features of the reference image into the video. On the one hand, by preserving the core functionality of self-attention and the pretrained prior of the video generation model, our method achieves the highest facial similarity and naturalness in identity-preserving video generation with only $\sim$1% additional parameters and 2000 training pairs, which is shown in Figure 2. On the other hand, our method does not alter the architecture of the main video generation model and thus can be used in a plug-and-play manner for other applications. As shown in Figure 1, our framework can be extended to various tasks, including subject-driven generation, video stylization, and face swapping, all guaranteeing identity consistency. Additionally, benefit from integrating compatibility with VACE [vace], our approach significantly enhances facial similarity in pose-guided video generation. Our main contributions can be summarized as follows:
@@ 2 Related Work
Current video generation models are predominantly built on diffusion frameworks [NEURIPS2020_4c5bcfec], with an evolution in architecture from U-Net-based designs [blattmann2023stablevideodiffusionscaling] to DiT-based approaches [kong2024hunyuanvideo, wan2.1, ma2025latte, DiT, wang2023laviehighqualityvideogeneration, VideoCrafter2, T2V-Turbo]. In the era of U-Net-based diffusion models, text-to-image (T2I) frameworks [rombach2021highresolution, podell2023sdxl] were extended to video generation by introducing 3D convolutions and temporal attention  [blattmann2023stablevideodiffusionscaling]. AnimateDiff [guo2023animatediff] further advanced this direction by reusing pre-trained text-to-image model weights to leverage their strong spatial generation capabilities by adding temporal layers. Latte [ma2025latte] introduced a spatial-temporal separation mechanism, assigning distinct DiT blocks to process spatial and temporal information independently. This approach was later replaced by 3D full attention mechanisms, which offered more integrated processing. CogVideoX [yang2024cogvideox] and HunyuanVideo [kong2024hunyuanvideo] combined 3D-VAE  [yu2023language] with MM-DiT [esser2024scaling] to enhance video generation capabilities. WAN2.1 [wan2.1] employs a 3D-VAE and adopts a DiT backbone for denoising, injecting semantic information into the diffusion process through cross-attention.
@@
Prior to the advent of zero-shot identity-preserving algorithms, generating content with consistent identity typically relied on case-by-case fine-tuning.[Dreambooth, Lora, textual_inversion, multi-concept, dreamvideo, customvideo, motionbooth], In contrast, training-free identity-preserving image generation methods[ye2023ip, shi2024instantbooth, guo2025pulid, wang2024instantid, he2024uniportrait, sara2025ipcompose, qian2024omni, han2024face, mou2025dreamounifiedframeworkimage] offer a zero-shot personalization solution by integrating identity features into pre-trained foundation models. These methods typically introduce parameterized plug-in modules or adapters to adjust and inject identity features into the generation process. A popular solution is to use a Face Encoder to extract face embeddings and inject them into the generation process via cross-attention. For example, InstantID[wang2024instantid] and PuLID[guo2025pulid] can generate high-quality, identity-preserving images.
In the field of identity-preserving video generation. Early methods commonly rely on explicit face encoders for facial feature extraction to generate videos with identity-preserving. ID-animator [ID-Animator] leverages a pre-trained text-to-video diffusion model in conjunction with a lightweight face adapter to encode ID-relevant embeddings from adaptable facial latent queries. ConsistID [consistid] aims to maintain identity consistency through frequency decomposition in diffusion transformer. Phantom [liu2025phantom] can also preserve identity consistency in the human domain as a unified subject-consistent video generation framework. HunyuanCustom [hu2025hunyuancustom] is a multi-modal customized video generation framework that emphasizes identity consistency while supporting diverse input modalities. By introducing advanced condition injection mechanisms and identity-preserving strategies, it achieves excellent performance in high-quality video generation. They employed a full fine-tuning for the diffusion transformer, requiring huge training resources.
@@ 3.1 Overall Framework
To extract facial features, early methods [ID-Animator, consistid] rely on explicit face encoders, which lack flexibility and often fail to preserve fine facial details critical for high-quality reconstruction. In contrast, we propose using the pre-trained VAE from the video generation model. This strategy maps the conditional image directly into the same latent space as the video and allows us to naturally take advantage of the built-in ability of the pre-trained video generation model to extract rich facial features.
The overall framework is illustrated in Figure 3. We use Wan2.1-14B-T2V [wan2.1] as the video generation base model, which adopts a Diffusion Transformer (DiT) architecture. Given a reference image containing a face, we first encode it into the latent space using the pre-trained VAE encoder. The image latents undergo the same patchification and encoding procedures as the video latents. Then, the image tokens are concatenated with video tokens along the sequence dimension and processed jointly through successive blocks. Finally, image tokens are discarded at the final layer.
[…]
@@ 3.3 Conditional Position Mapping
To effectively differentiate image and video tokens in the restricted self-attention, we use a specialized conditional position mapping strategy. Specifically, we employ 3D Rotary Positional Embedding (RoPE) [RoPE], where all tokens associated with the reference image are assigned a distinct and dedicated coordinate space. This ensures clear separation and facilitates precise modeling of interactions between the reference image and video tokens.
@@
This non-overlapping spatial allocation achieves two main goals through geometric separation. By separating the reference tokens from the video coordinate grid, the design naturally reduces false spatial correlations and better preserves the backbone’s pretrained positional prior. Compared with Shared Position Mapping, our Conditional Position Mapping generates more reliable video with more stable scene layouts, which is shown in Figure 6. At the same time, this separate coordinate system maintains the reference image’s semantic meaning by keeping it as a global identity prior. Consequently, the model is guided to focus on extracting overall semantic features from the reference tokens, rather than treating them as spatially localized patterns that need to align positionally with the video content.
@@ 3.4 Dataset Collection and Processing
We construct a human-centric video dataset containing 2,000 high-resolution sequences from publicly available sources. The dataset guarantees a diverse and comprehensive representation, comprising various ethnic groups, age ranges, gender identities, and a wide array of actions. Using the VILA [lin2024vila] multimodal captioning framework, we automatically generate dense textual annotations for each video, establishing strong text-video alignment.
@@
    Faces are detected and cropped using RetinaFace[Deng2020CVPR].
@@
    BiSeNet [BiseNet] is used for face parsing, and the background is replaced with a solid white color to prevent any leakage of background information.
@@
ID-Animator [ID-Animator] 0.316 3.211 16.677
SkyReels-A2-P14B [SkyReels] 0.546 3.411 19.110
EchoVideo [EchoVideo] 0.487 3.456 19.263
ConcatID-CogVideoX [Concat-ID] 0.439 3.372 19.359
ConcatID-WAN [Concat-ID] 0.501 3.650 19.671
Hunyuan-Custom [HunyuanCustom] 0.622 3.367 19.853
ConsistID [consistid] 0.432 3.233 20.552
VACE-P1.3B [vace] 0.180 3.567 20.591
VACE-1.3B [vace] 0.223 3.611 20.527
VACE-14B [vace] 0.647 3.728 19.520
Phantom-1.3B [liu2025phantom] 0.440 3.567 20.364
Phantom-14B [liu2025phantom] 0.519 3.828 20.476
@@ 4.1 Implementation Details
Hunyuan-Custom[HunyuanCustom] 3.34 2.92
VACE-14B [vace] 3.00 3.07
Phantom-14B [liu2025phantom] 2.37 2.92
ConsistID [consistid] 2.25 2.46
@@ 4.2 Quantitative Analysis
To evaluate identity preservation and visual quality, we use the two most important and heavily weighted evaluation metrics from the OpenS2V benchmark [yuan2025opens2vnexusdetailedbenchmarkmillionscale]: facial similarity and naturalness. To evaluate the relevance between generated video and textual description, we use X-CLIP [xclip], a pre-trained video-text multimodal model. Results are shown in Table 1 and Figure 8. As shown in Figure 9, despite being trained with only $\sim$1% additional parameters and just 2000 training pairs, our model also demonstrates strong generalization across individuals of diverse ethnicities and age groups.
@@ 4.4 Plug-and-Play to Other Applications
The proposed conditional image branch is designed based on the LoRA module, which ensures inherent compatibility with other DiT-based models. To validate this, we conducted experiments on the pose-guided video generation task using the VACE framework [vace]. As illustrated in Figure 11, integrating our method significantly improves the facial identity similarity in the generated videos. This not only demonstrates the plug-and-play nature of our approach but also highlights its robustness in preserving identity consistency.
@@
Our framework is also capable of video face swapping, which can be achieved via zero-shot inpainting [inpainting]. Figure 12 shows that our method not only achieves high-quality facial identity transfer but also maintains strong temporal consistency across frames, resulting in coherent and high-quality videos.
@@ 5 Conclusion
Stand-In: A Lightweight and Plug-and-Play Identity Control
for Video Generation
     Supplemental Material
6 Pseudocode
To more clearly and systematically present the core inference mechanism of the Stand-In method, we formalize its complete inference procedure in Algorithm 1. This pseudocode illustrates how the model jointly leverages the reference image during the diffusion process to preserve identity consistency, thereby enabling high-quality, identity-preserving video generation.
Algorithm 1 Inference Process of Stand-In
Input: Reference image tokens $T_{I}$, initial video latents $T_{V_{N}}$, list of timesteps $S=\{s_{N},s_{N-1},...,s_{1}\}$, DiT model $\mathcal{M}$.
Parameter: Zero timestep for the reference image, $s_{ref}=0$; Positional Encodings $p_{I},p_{V}$.
Output: Denoised video latents $T_{V_{0}}$.
1: Initialize an empty cache for each DiT block: $C_{k},C_{v}\leftarrow\{\},\{\}$
[…]
3: Let $is\_first\_step\leftarrow\text{true}$
4:for $s$ in $S$ do
5:  for each block $\mathcal{B}$ in model $\mathcal{M}$ do
6:   if is_first_stepis\_first\_step then
7:    Compute $(Q_{I},K_{I},V_{I})\leftarrow\mathcal{B}_{\text{before\_self-attn}}(T_{I},s_{ref})$
8:    Store $K_{I},V_{I}$ in cache: $C_{k}[\mathcal{B}]\leftarrow K_{I}$, $C_{v}[\mathcal{B}]\leftarrow V_{I}$
9:   else
10:    Retrieve from cache: $K_{I},V_{I}\leftarrow C_{k}[\mathcal{B}],C_{v}[\mathcal{B}]$
11:   end if
12:   Compute $(Q_{V},K_{V},V_{V})\leftarrow\mathcal{B}_{\text{before\_self-attn}}(T_{V},s)$
13:   // Self-attention start
14:   $(Q_{V}^{\prime},K_{V}^{\prime})\leftarrow\textnormal{{ApplyRoPE}}((Q_{V},K_{V}),p_{V})$
15:   $(Q_{I}^{\prime},K_{I}^{\prime})\leftarrow\textnormal{{ApplyRoPE}}((Q_{I},K_{I}),p_{I})$
16:   $K^{\prime}_{\text{cat}}\leftarrow\text{Concat}(K_{V}^{\prime},K_{I}^{\prime})$, $V_{\text{cat}}\leftarrow\text{Concat}(V_{V},V_{I})$
17:   $\text{Out}_{I}\leftarrow\text{Attention}(Q_{I}^{\prime},K_{I}^{\prime},V_{I})$
18:   $\text{Out}_{V}\leftarrow\text{Attention}(Q_{V}^{\prime},K^{\prime}_{\text{cat}},V_{\text{cat}})$
19:   // Self-attention end
20:   $(T_{V},T_{I})\leftarrow\mathcal{B}_{\text{after\_self-attn}}(\text{Out}_{V},\text{Out}_{I})$
21:  end for
22:  $is\_first\_step\leftarrow\text{false}$
23:  Predict denoised $T_{V}$ using scheduler and update for next step.
24:end for
25:return $T_{V_{0}}$
7 More Comparison Results
We provide more comparisons of our model with other identity-preserving video generation methods in Figures 14-17.
Figure 14: Visual comparison with other cutting-edge methods. Our method, Stand-In, demonstrates superior identity fidelity and detail preservation (1/4).
Figure 15: Visual comparison with other cutting-edge methods. Our method, Stand-In, demonstrates superior identity fidelity and detail preservation (2/4).
Figure 16: Visual comparison with other cutting-edge methods. Our method, Stand-In, demonstrates superior identity fidelity and detail preservation (3/4).
Figure 17: Visual comparison with other cutting-edge methods. Our method, Stand-In, demonstrates superior identity fidelity and detail preservation (4/4).
8 More Ablation Study Results
To further validate the effectiveness of our proposed Restricted Self-Attention (RSA) and Conditional Position Mapping (CPM), we provide additional qualitative comparisons in Figure 18 . The ablation focuses on two alternative configurations: (1) replacing RSA with vanilla self-attention (VSA), and (2) replacing CPM with Shared Position Mapping (SPM).
When RSA is replaced with Vanilla Self-Attention (VSA), the model suffers from severe identity drift. This confirms that unconstrained bidirectional attention causes the reference-image tokens to be contaminated by dynamic video content, preventing stable identity extraction. Similarly, when CPM is replaced with Shared Position Mapping (CPM → SPM), sharing the positional space results in clearly distorted facial proportions (e.g., enlarged face in the second example), revealing that the pretrained positional prior is severely disrupted. In contrast, our full model consistently maintains both accurate identity and natural facial geometry for both cases, demonstrating that RSA and CPM are jointly crucial for stable, identity-preserving video generation.
Figure 18: Visual results of the ablation study.
9 More Generalization Results
To further evaluate the robustness and generalization ability of our approach, Figure 19 showcases identity-preserving video generation results for a diverse set of subjects. Despite significant variations in age, ethnicity, skin tone, and facial geometry, our method consistently reconstructs key identity traits and maintains stable facial structure across frames. The generated videos exhibit natural motion, coherent expressions, and high perceptual fidelity. The text prompts from top to bottom are:
(1) A white man stands in front of his desk, facing the camera while speaking and gesturing to emphasize his points. He then picks up a pen and writes a few lines on a sheet of paper before looking back up with a confident smile, as if showing his progress to the viewer. The lighting is soft, giving him a natural and self-assured presence.
(2) An elderly Asian woman sits on a wooden balcony chair, smiling gently at the camera. She takes a small sip from her teacup, then looks up with a warm smile and nods. She reaches to adjust the shawl draped over her shoulders and closes her eyes slightly to enjoy the sunlight. Her expression is peaceful and kind, giving a sense of warmth and comfort.
(3) An elderly Black man sits on his home balcony, looking directly at the camera while holding a warm cup of tea with both hands. A gentle breeze brushes past as he softly blows across the tea’s surface and looks up with a warm smile. He then raises his cup and nods toward the camera in a calm and friendly gesture.
(4) A white woman faces the camera, gently sweeping a makeup brush across her face. She then offers a radiant smile, tilting her head slightly to show her makeup before returning to the front and giving a playful wink. Her movements are graceful and natural, complemented by soft lighting.
(5) An Asian man faces the camera, lightly adjusting the collar of his shirt before running his hand through his hair with a smile. He then raises a cup of tea and makes a small “cheers” gesture toward the camera. His expression is natural and confident, and his movements feel relaxed and smooth.
(6) A young Black woman smiles at the camera while adjusting her curly hair. She gently shakes a coffee cup toward the lens, as if sharing her morning energy with the viewer. She blinks playfully, makes an “OK” gesture, and smiles as sunlight from the window falls softly across her face, giving her a confident and natural expression.
(7) A young white boy smiles at the camera, giving a small wave. He glances down with a shy laugh before looking up again with a bright and innocent expression. Sunlight illuminates his face, making him look natural and adorable.
(8) A young Asian girl faces the camera and makes a heart shape with her hands. She then picks up a glass of juice, takes a sip, and smiles as she wipes the corner of her mouth. A breeze lifts her hair as she playfully brushes it aside, her eyes bright and expressive.
(9) A young Black girl smiles brightly at the camera. She waves hello softly, then forms a heart shape with her hands. Her eyes are bright, her smile sweet, and the sunlight behind her creates a warm and gentle atmosphere.
Figure 19: Given a single reference image (left), our method generates videos (right) that maintain strong identity consistency across diverse ethnicities and age groups, despite being trained with only $\sim$1% additional parameters and just 2000 training pairs.
10 Complete Video Generation Prompts
In this section, we present the complete prompts for Figures in the main paper.
Figure 8: Comparison on identity-preserving video generation
Left Panel:
The video features a man standing at an easel, focused intently as his brush dances across the canvas. His expression is one of deep concentration, with a hint of satisfaction as each brushstroke adds color and form. He wears a paint-splattered apron, and his hands move with confident precision. The setting, filled with scattered art supplies, open paint tubes, and unfinished sketches pinned to the wall, suggests an artist’s studio. A large window on one side allows sunlight to stream in, casting a soft glow across the room and illuminating the colors on his canvas. The atmosphere is creative and inspired, with the man’s intense focus and the lively colors on the canvas indicating a moment of artistic passion and expression.
Right Panel:
The video features a man with dark-haired hair, wearing a blue tank top and holding a pink tank top on a hanger. he appears to be in a clothing store or a similar retail environment, as there are racks of clothes visible in the background. The man is speaking to the camera, possibly providing a review or discussing the tank top he is holding. He has colorful bracelets on his wrist and is wearing a necklace with multiple beads. his expression suggests he is engaged in a conversation or presentation. The setting seems to be indoors, with artificial lighting illuminating the scene.
Figure 10: Our results on subjects other than real-person
Left Panel:
Shot in a medium shot of a brightly lit room, a girl, approximately seven or eight years old, stands in the center. She has long black hair and wears a light blue dress, her expression focused and gentle. Holding a doll in both hands, she presents her beloved toy to the camera. As the camera slowly zooms in, the details of her face are clearly visible: the soft fabric, the delicate stitching, and the slightly upturned corners of her mouth are all captured. The entire scene is filled with childlike innocence and warmth.
Right Panel:
The video features an anime girl standing on a busy street, surrounded by a hurried crowd. The buildings and shops in the background create a classic cityscape. The girl smiles as she puts her headphones on, her movements smooth and natural. Her expression is playful and relaxed, as if she’s about to immerse herself in her favorite music. The camera focuses on her face, capturing her joyful expression and vibrant energy. The background is slightly blurred, emphasizing the contrast between her and her surroundings, creating a sense of relaxed urban living.
Figure 12: Application of our model in video face swapping
Top Panel:
A woman is seated comfortably at a desk, facing the camera directly as if engaged in a video call with a friend or family member. Her gaze is intent and warm, complemented by a natural smile. The background consists of a personalized space, with photographs and a world map adorning the wall, thereby communicating a feeling of intimate and contemporary connection.
Bottom Panel:
A man is seated in a comfortable armchair, holding a thick, aged book on his lap. Drawn by the camera’s presence, he looks up, his face bearing a gentle and curious smile as if mildly interrupted. Soft, lateral lighting illuminates the scene. He is surrounded by floor-to-ceiling bookshelves, creating a tranquil atmosphere enveloped by knowledge and narrative.
11 More Results
Figure 20: We present more visual results of Stand-In (1/2).
Figure 21: We present more visual results of Stand-In (2/2).

### v4

@@ 1 Introduction
With the rapid advancement of diffusion models [11, 30, 28], video generation [52, 29, 17, 12] has become a pivotal aspect of generative AI. Among its diverse applications, identity-preserving video generation holds profound significance. The goal of this task is to generate high-quality videos that consistently maintain the identity of a given reference image containing a face. It has widespread utility across film, advertising and gaming industries, etc.
Existing methods can be roughly classified into two categories. Early methods [10, 50] use an explicit face encoder for identity feature extraction, while recent methods [14, 21] fully train the diffusion transformer. However, face-encoder-based methods lack flexibility and struggle to capture fine facial details essential for high-quality video generation. The full-parameter training methods require huge training resources and lack compatibility with other applications. Achieving robust identity preservation in a lightweight way remains critical yet challenging.
To overcome these limitations, we leverage the pre-trained VAE from the video generation model itself, enabling the conditional image to be mapped directly into the same latent space as the video. This approach naturally utilizes the model’s inherent capabilities to extract rich and detailed facial features, offering a more integrated and effective solution. Specifically, we employ restricted self-attention with conditional position mapping to merge the features of the reference image into the video. On the one hand, by preserving the core functionality of self-attention and the pretrained prior of the video generation model, our method achieves the highest facial similarity and naturalness in identity-preserving video generation with only $\sim$1% additional parameters and 2000 training pairs, which is shown in Figure 2. On the other hand, our method does not alter the architecture of the main video generation model and thus can be used in a plug-and-play manner for other applications. As shown in Figure 1, our framework can be extended to various tasks, including subject-driven generation, video stylization, and face swapping, all guaranteeing identity consistency. Additionally, benefit from integrating compatibility with VACE [16], our approach significantly enhances facial similarity in pose-guided video generation. Our main contributions can be summarized as follows:
@@ 2 Related Work
Current video generation models are predominantly built on diffusion frameworks [11], with an evolution in architecture from U-Net-based designs [1] to DiT-based approaches [17, 38, 25, 24, 40, 2, 19]. In the era of U-Net-based diffusion models, text-to-image (T2I) frameworks [32, 30] were extended to video generation by introducing 3D convolutions and temporal attention  [1]. AnimateDiff [6] further advanced this direction by reusing pre-trained text-to-image model weights to leverage their strong spatial generation capabilities by adding temporal layers. Latte [25] introduced a spatial-temporal separation mechanism, assigning distinct DiT blocks to process spatial and temporal information independently. This approach was later replaced by 3D full attention mechanisms, which offered more integrated processing. CogVideoX [45] and HunyuanVideo [17] combined 3D-VAE  [48] with MM-DiT [4] to enhance video generation capabilities. WAN2.1 [38] employs a 3D-VAE and adopts a DiT backbone for denoising, injecting semantic information into the diffusion process through cross-attention.
@@
Prior to the advent of zero-shot identity-preserving algorithms, generating content with consistent identity typically relied on case-by-case fine-tuning.[33, 13, 5, 18, 43, 41, 44], In contrast, training-free identity-preserving image generation methods[46, 35, 7, 39, 9, 34, 31, 8, 27] offer a zero-shot personalization solution by integrating identity features into pre-trained foundation models. These methods typically introduce parameterized plug-in modules or adapters to adjust and inject identity features into the generation process. A popular solution is to use a Face Encoder to extract face embeddings and inject them into the generation process via cross-attention. For example, InstantID[39] and PuLID[7] can generate high-quality, identity-preserving images.
In the field of identity-preserving video generation [23, 51], early methods commonly rely on explicit face encoders for facial feature extraction to generate videos with identity-preserving. ID-animator [10] leverages a pre-trained text-to-video diffusion model in conjunction with a lightweight face adapter to encode ID-relevant embeddings from adaptable facial latent queries. ConsistID [50] aims to maintain identity consistency through frequency decomposition in diffusion transformer. Phantom [21] can also preserve identity consistency in the human domain as a unified subject-consistent video generation framework. HunyuanCustom [14] is a multi-modal customized video generation framework that emphasizes identity consistency while supporting diverse input modalities. By introducing advanced condition injection mechanisms and identity-preserving strategies, it achieves excellent performance in high-quality video generation. They employed a full fine-tuning for the diffusion transformer, requiring huge training resources.
@@ 3.1 Overall Framework
To extract facial features, early methods [10, 50] rely on explicit face encoders, which lack flexibility and often fail to preserve fine facial details critical for high-quality reconstruction. In contrast, we propose using the pre-trained VAE from the video generation model. This strategy maps the conditional image directly into the same latent space as the video and allows us to naturally take advantage of the built-in ability of the pre-trained video generation model to extract rich facial features.
The overall framework is illustrated in Figure 3. We use Wan2.1-14B-T2V [38] as the video generation base model, which adopts a Diffusion Transformer (DiT) architecture. Given a reference image containing a face, we first encode it into the latent space using the pre-trained VAE encoder. The image latents undergo the same patchification and encoding procedures as the video latents. Then, the image tokens are concatenated with video tokens along the sequence dimension and processed jointly through successive blocks. Finally, image tokens are discarded at the final layer.
To preserve the static nature of the reference image, which serves as a conditioning input rather than undergoing denoising, we maintain its temporal invariance. This is done by fixing its timestep to zero, i.e.  $s_{ref}=0$, where $s$ denotes the denoising timestep in diffusion. Now that we have encoded the conditional image into the same feature space as the video, the next challenge is: How can the video features effectively refer to the image information in a way that is lightweight and easy to learn?
@@ 3.3 Conditional Position Mapping
To effectively differentiate image and video tokens in the restricted self-attention, we use a specialized conditional position mapping strategy. Specifically, we employ 3D Rotary Positional Embedding (RoPE) [36], where all tokens associated with the reference image are assigned a distinct and dedicated coordinate space. This ensures clear separation and facilitates precise modeling of interactions between the reference image and video tokens.
@@
This non-overlapping spatial allocation achieves two main goals through geometric separation. By separating the reference tokens from the video coordinate grid, the design naturally reduces false spatial correlations and better preserves the backbone’s pretrained positional prior. Compared with Shared Position Mapping, our Conditional Position Mapping generates more reliable video with more stable proportions, which is shown in Figure 6. At the same time, this separate coordinate system maintains the reference image’s semantic meaning by keeping it as a global identity prior. Consequently, the model is guided to focus on extracting overall semantic features from the reference tokens, rather than treating them as spatially localized patterns that need to align positionally with the video content.
@@ 3.4 Dataset Collection and Processing
We construct a human-centric video dataset containing 2,000 high-resolution sequences from publicly available sources. The dataset guarantees a diverse and comprehensive representation, comprising various ethnic groups, age ranges, gender identities, and a wide array of actions. Using the VILA [20] multimodal captioning framework, we automatically generate dense textual annotations for each video, establishing strong text-video alignment.
@@
    Faces are detected and cropped using RetinaFace[3].
@@
    BiSeNet [47] is used for face parsing, and the background is replaced with a solid white color to prevent any leakage of background information.
@@
ID-Animator [10] 0.316 3.211 16.677
SkyReels-A2-P14B [37] 0.546 3.411 19.110
EchoVideo [42] 0.487 3.456 19.263
ConcatID-CogVideoX [53] 0.439 3.372 19.359
ConcatID-WAN [53] 0.501 3.650 19.671
Hunyuan-Custom [15] 0.622 3.367 19.853
ConsistID [50] 0.432 3.233 20.552
VACE-P1.3B [16] 0.180 3.567 20.591
VACE-1.3B [16] 0.223 3.611 20.527
VACE-14B [16] 0.647 3.728 19.520
Phantom-1.3B [21] 0.440 3.567 20.364
Phantom-14B [21] 0.519 3.828 20.476
@@ 4.1 Implementation Details
Hunyuan-Custom[15] 3.34 2.92
VACE-14B [16] 3.00 3.07
Phantom-14B [21] 2.37 2.92
ConsistID [50] 2.25 2.46
@@ 4.2 Quantitative Analysis
To evaluate identity preservation and visual quality, we use the two most important and heavily weighted evaluation metrics from the OpenS2V benchmark [49]: facial similarity and naturalness. To evaluate the relevance between generated video and textual description, we use X-CLIP [26], a pre-trained video-text multimodal model. Results are shown in Table 1 and Figure 8. As shown in Figure 9, despite being trained with only $\sim$1% additional parameters and just 2000 training pairs, our model also demonstrates strong generalization across individuals of diverse ethnicities and age groups.
@@ 4.4 Plug-and-Play to Other Applications
The proposed conditional image branch is designed based on the LoRA module, which ensures inherent compatibility with other DiT-based models. To validate this, we conducted experiments on the pose-guided video generation task using the VACE framework [16]. As illustrated in Figure 11, integrating our method significantly improves the facial identity similarity in the generated videos. This not only demonstrates the plug-and-play nature of our approach but also highlights its robustness in preserving identity consistency.
@@
Our framework is also capable of video face swapping, which can be achieved via zero-shot inpainting [22]. Figure 12 shows that our method not only achieves high-quality facial identity transfer but also maintains strong temporal consistency across frames, resulting in coherent and high-quality videos.
@@ 5 Conclusion
6 Acknowledgment
This work was supported in part by the National Natural Science Foundation of China (62306153, 62225604), Tianjin Natural Science Foundation Project (25ZXRGGX00290, 24JCJQJC00020, 25JCQNJC01390), the Young Elite Scientists Sponsorship Program by CAST (YESS20240686), the Fundamental Research Funds for the Central Universities (Nankai University, 63253223, 63253219), and Shenzhen Science and Technology Program (JCYJ20240813114237048). The computational devices is supported by the Supercomputing Center of Nankai University (NKSC).